OpenStudy (anonymous):

find the distance between the points with these coordinates. Use the distance formula when helpful or necessary. points (0, 0) and (6, -8) Distance:______

Parth (parthkohli):

Use the distance formula :)

Parth (parthkohli):

\(\Large \color{Orange}{\Rightarrow \sqrt{(-8 - 0)^2 + (6 - 0)^2} }\)

OpenStudy (anonymous):

I did but I I got square root of 28, and it still wrong..

OpenStudy (anonymous):

\[\large \sqrt {(-8-0)^2+(6-0)^2} \Rightarrow \sqrt {64+36}= \]